# Stacky Batyrev–Manin conjecture and modular curves

### v1

@@ 1.1. Batyrev–Manin conjecture and stacks
Let us first define this cone. A stacky curve on $\mathcal{X}_{\overline{F}}$ is a representable $\overline{F}$-morphism $f:\mathcal{C}\to\mathcal{X}_{\overline{F}}$ with $\mathcal{C}$ stacky curve (which is a smooth separated irreducible stack of dimension $1$). By [GS17, Theorem 1] we can assume that the automorphism group schemes of points of $\mathcal{C}$ are $\mu_{k}$ for some $k$. At each stacky point $z$ of $\mathcal{C}$, we get an associated sector $\psi(z)$ of $\mathcal{X}$, it is the one defined by the pair $(f(z),\operatorname{Aut}(z)\hookrightarrow\operatorname{Aut}(f(z))).$ For $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines an intersection number $(f:\mathcal{C}\to\mathcal{X}_{\overline{F}},(L,x))$ by setting it to be the sum of $\deg(f^{*}L)$ and $\sum_{z\in S(\mathcal{C})}x_{\psi(z)}$, where $S(\mathcal{C})$ is the set of stacky points of $\mathcal{C}$ and $x_{0}$ is interpreted as 0. A family of movable stacky curves on $\mathcal{X}_{\overline{F}}$ is a family of stacky curves on $\mathcal{X}_{\overline{F}}$ which dominates $\mathcal{X}_{\overline{F}}$. One defines an intersection number between a family of movable curves and $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, as the generic value of the intersection of the members of the family with $(L,x)$ [DY24, Definition 8.7]. Motivated by the fact that the effective cone is the dual of the cone of movable curves [BDPP13], one defines $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})$ to be the subset of $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$ which intersect non-negatively every movable family of stacky curves on $\mathcal{X}_{\overline{F}}$. If a height is defined by $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines
@@
Finally, we need one more definition before stating the main conjecture: given a stack $\mathcal{Y}$ and a ring $R$, denote by $\mathcal{Y}\langle R\rangle$ the set of isomorphism classes of $R$-points of $\mathcal{Y}$.
@@ 1.2. Results
The main theorem of this paper proves the Conjecture 1.1 on several modular curves. Recall that given a stack $\mathcal{X}$ defined over a number field $F$, $\mathcal{X}\langle F\rangle$ is the set of isomorphism classes of $F$-points of $\mathcal{X}$. For each $N\in\mathbb{Z}_{>0}$, let $\mathscr{X}_{0}(N)$ be the compactified Deligne–Rapoport modular stack of elliptic curves endowed with a cyclic rational $N$-isogeny over $F$ (see Definition 3.2 for the precise definition). Consider
@@
Then the coarse moduli space $X_{0}(N)$ of $\mathscr{X}_{0}(N)$ is geometrically rationally connected if and only if it has genus 0, that is if and only if $N\in\mathfrak{N}_{0}$ (c.f. [Maz78]). So we assume that $N\in\mathfrak{N}_{0}$. The $F$-points of these stacks of bounded naive height have been counted by the papers [BS24, Mol23, MV23, AHPP25]. When $N=1$, by the naive height on $\mathscr{X}_{0}(1)\langle F\rangle\cong\mathcal{P}(4,6)\langle F\rangle\to\mathbb{R}_{\geq 0},$ we mean the height function from Definition 3.9. When $F=\mathbb{Q}$, the $\mathbb{Q}$-elliptic curve given by $Y^{2}=X^{3}+eX+f,$ where $e,f\in\mathbb{Z}$ satisfy that for all $p$ prime, one has $p^{4}\nmid e$ or $p^{6}\nmid f$, the naive height is given by $\max(|e^{3}|,|f^{2}|)$. (Note that some sources use a different normalization, e. g. in [AHPP25], it is given by $\max(4|e^{3}|,27|f^{2}|),$ but it has no effect on $a$- and $b$-invariants.) By naive height on $\mathscr{X}_{0}(N)\langle F\rangle$, we mean the composite function $H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathscr{X}_{0}(1)\langle F\rangle\to\mathbb{R}_{>0}$, where $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ forgets the level structure. By Lemma 3.4, one has that $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
@@
Let $N\in\mathfrak{N}_{0}$ as above and let $M_{\operatorname{naive}}\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ be the element corresponding to the naive height $H$. Assume that $F=\mathbb{Q}$. Then Conjecture 1.1 holds for $\mathscr{X}_{0}(N)$ with the respect to the height $H$.
@@
For each $N$ and $M_{\operatorname{naive}}$ from Theorem 1.2, one has the following values of $a$- and $b$-invariants:
@@
To prove Theorem 1.3, we first give a concrete description of $\mathscr{X}_{0}(N)$ as a square root stack over a stacky curve. Recall that given a line bundle $L$ on a stack $\mathcal{X}$, $n$-th root stack of $L$, $\sqrt[n]{L/\mathcal{X}}$, is a stack defined in [Cad07], endowed with a morphism to $\mathcal{X}$ such that the pullback of $L$ is $n$-th power in the Picard group of $\sqrt[n]{L/\mathcal{X}}$.
The square root stacks are $\mu_{2}$-gerbes: they are the trivial $\mu_{2}$-gerbes if and only if $L$ is a square in the Picard group $\operatorname{Pic}(\mathcal{X})$. The square root stacks correspond to $\operatorname{Pic}(X)/2\operatorname{Pic}(X)$. In [AOV08], one describes a notion of rigidification, which essentially kills a subgroup scheme of generic stabilizer group. Hence if $\mathcal{X}$ is a stacky curve and $L$ line bundle, the generic stabilizer is $\mu_{2}$ and the rigidification of $\sqrt[2]{L/\mathcal{X}}$ is just $\mathcal{X}$. We denote the rigidification of $\mathcal{Y}$ by $\mathcal{Y}^{\operatorname{rig}}$.
Similarly, to notion of root stack of a line bundle, there is a notion of $(n_{1},\dots,n_{r})$-th root stack of an $r$-tuple of Cartier divisors $C_{1},\dots,C_{r}$ on an algebraic stack $\mathcal{X}$. When the dimension of $\mathcal{X}$ is equal to $1$, this the stack $\sqrt[n_{1},\dots,n_{r}]{C_{1},\dots,C_{r}/\mathcal{X}}\to X$ can be thought as adding stackiness $\mu_{n_{i}}$ along $C_{i}$. Let $K_{6}$ and $K_{4}$ be the divisors $X^{2}+3Y^{2}=0$ and $X^{2}+Y^{2}=0$, respectively, on $\mathbb{P}^{1}$. Let $E_{0},E_{1728}$ be the divisors corresponding to $X=0$ and $X=1728Y$, respectively.
We have a canonical representable morphism $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)=\mathscr{X}_{0}(1)$ and an induced representable morphism $J_{N}^{\operatorname{rig}}:\mathscr{X}_{0}(N)^{\operatorname{rig}}\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. Let $\Lambda$ be the Hodge line bundle on $\mathscr{X}_{0}(1)$. There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$ such that $\Lambda^{\otimes 2}$ is isomorphic to the pullback of $\Lambda^{\operatorname{rig}}$ under the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. We set $\Lambda_{N}:=J_{N}^{*}\Lambda$ and $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}\Lambda^{\operatorname{rig}}$.
@@ 1.4. Organization
In Section 2, we recall cyclotomic stacks, stacky Proj construction, root stacks, and sectors of DM stacks. Section 3 starts with the definition and geometric properties of modular curves $\mathscr{X}_{0}(N)$. After that, the naive height of $\mathscr{X}_{0}(N)$ is studied in Section 3.1, which concludes with the description of the element defining the naive height in $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ (Corollary 3.13). In Section 3.2, we consider the orbifold canonical divisor $K_{\mathscr{X}_{0}(N),\operatorname{orb}}$ and the anticanonical naive height. In Section 3.3, we classify adequate cases and then prove Theorem 1.3 for these cases. Section 4 starts with Section 4.1, where the local computations are discussed to understand stacky curves over $\mathscr{X}_{0}(N)$. Then definitions and properties of the orbifold pseudoeffective cone $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ are described in Section 4.2. In the remaining subsections of Section 4, Theorem 1.3 is proved for the non-adequate cases.
@@ 2.1. Cyclotomic stacks and the stacky Proj construction
Given a $\mathbb{Q}$-stack $\mathcal{X}$, [AHPP25, Lemma 2.7] explains necessary and sufficient conditions for $\mathcal{X}$ to be a stacky Proj, which involves definitions from [AH11, §2]; so let us summarize those definitions over $\operatorname{Spec}F$. A separated $F$-stack $\mathcal{X}$ of locally finitely presentation over $\operatorname{Spec}F$ is called a cyclotomic stack if for every geometric point $\overline{x}$ of $X$, the stabilizer group space $\underline{\operatorname{Aut}}_{\mathcal{X}}(\overline{x})$ (or $\underline{\operatorname{Aut}}(\overline{x})$ for short) is isomorphic to $(\mu_{n})_{\kappa(\overline{x})}$ for some $n\geq 1$. A line bundle $\mathcal{L}$ on a cyclotomic $F$-stack $\mathcal{X}$ is uniformizing if the associated classifying morphism $\mathcal{X}\to B\mathbb{G}_{m}$ is representable; equivalently, for every geometric point $\overline{x}$ of $\mathcal{X}$, $\underline{\operatorname{Aut}}(\overline{x})$-action on $\mathcal{L}_{\overline{x}}$ is faithful (c.f. [AH11, Proposition 2.3.10]). Finally, a uniformizing line bundle $\mathcal{L}$ on a proper cyclotomic $F$-stack $\mathcal{X}$ is polarizing if there exists $m\geq 1$ such that $\mathcal{L}^{m}\cong c^{*}M$, where $c:\mathcal{X}\to X$ is the coarse moduli map and $M$ is an ample line bundle on $X$. A mild generalization of [AHPP25, Lemma 2.7] is as follows:
@@ 2.2. Root stacks
Given an effective Cartier divisor $\mathcal{D}$ on $\mathcal{X}$, pick a global section $s_{\mathcal{D}}$ of $\mathcal{O}_{\mathcal{X}}(\mathcal{D})$ that cuts out $\mathcal{D}$. Then the $n^{\rm th}$ root stack of $\mathcal{D}$ on $\mathcal{X}$ is defined as the fiber product $\sqrt[n]{\mathcal{D}/\mathcal{X}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{X}\times_{[\mathbb{A}^{1}/B\mathbb{G}_{m}]}[\mathbb{A}^{1}/B\mathbb{G}_{m}]$ of $\mathcal{X}\to[\mathbb{A}^{1}/B\mathbb{G}_{m}]$ corresponding to $(\mathcal{L},s_{\mathcal{D}})$ and the $n^{\rm th}$-power morphism $\mathbb{A}^{1}/B\mathbb{G}_{m}]\to[\mathbb{A}^{1}/B\mathbb{G}_{m}]$; the isomorphism class of $\sqrt[n]{\mathcal{L}/\mathcal{X}}$ is independent of the choice of $s_{\mathcal{D}}$. Similarly, given a fixed choice $s_{\mathcal{D}}$, a $T$-object of $\sqrt[n]{\mathcal{D}/\mathcal{X}}$ is instead a quadruple $(x,M,u,\varphi)$ where $u\in H^{0}(M)$ and $\varphi(u^{n})=x^{*}s_{\mathcal{D}}$. Finally, when $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ are mutually disjoint effective Cartier divisors of $\mathcal{X}$, then $\sqrt[n_{1},\dotsc,n_{m}]{\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}/\mathcal{X}}$ means iterative root stack $\sqrt[n_{1}]{\mathcal{D}_{1}/\sqrt[n_{2}]{\mathcal{D}_{2}/\dotsb/\sqrt[n_{m}]{\mathcal{D}_{m}/\mathcal{X}}}}$; note that the isomorphism class of this stack remains the same under permutation of $(n_{1},\mathcal{D}_{1}),\dotsc,(n_{m},\mathcal{D}_{m})$. Because our stacks of interests are DM stacks of dimension one, the mutually disjoint condition of divisors $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ is not restrictive.
@@
Observe that the rigidification morphism $\pi\colon\mathcal{X}\to\mathcal{X}^{\operatorname{rig}}$ is a $\mu_{n}$-gerbe by definition. Because $\pi^{*}\mathcal{L}^{\operatorname{rig}}\cong\mathcal{L}^{n}$, this isomorphism induces a morphism $h\colon\mathcal{X}\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$ of $\mu_{n}$-gerbes over $\mathcal{X}^{\operatorname{rig}}$. Then for every smooth $F$-covering $U\to\mathcal{X}^{\operatorname{rig}}$ by an $F$-scheme $U$, the pullback $h_{U}\colon\mathcal{X}\times_{\mathcal{X}^{\operatorname{rig}}}U\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}\times_{\mathcal{X}^{\operatorname{rig}}}U$ is an isomorphism by [Ols16, Lemma 12.2.4], hence $h$ is also an isomorphism by descent. ∎

### v2

# The Stacky Batyrev–Manin conjecture and modular curves

@@ 1.1. Batyrev–Manin conjecture and stacks
Let us first define this cone. A stacky curve on $\mathcal{X}_{\overline{F}}$ is a representable $\overline{F}$-morphism $f:\mathcal{C}\to\mathcal{X}_{\overline{F}}$ with $\mathcal{C}$ stacky curve (which is a smooth separated irreducible stack of dimension $1$). By [GS17, Theorem 1] we can assume that the automorphism group schemes of points of $\mathcal{C}$ are $\mu_{k}$ for some $k$. At each stacky point $z$ of $\mathcal{C}$, we get an associated sector $\psi(z)$ of $\mathcal{X}$ which is defined by the pair $(f(z),\operatorname{Aut}(z)\hookrightarrow\operatorname{Aut}(f(z))).$ For $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines an intersection number $(f:\mathcal{C}\to\mathcal{X}_{\overline{F}},(L,x))$ by setting it to be the sum of $\deg(f^{*}L)$ and $\sum_{z\in S(\mathcal{C})}x_{\psi(z)}$, where $S(\mathcal{C})$ is the set of stacky points of $\mathcal{C}$ and $x_{0}$ is interpreted as 0. A family of movable stacky curves on $\mathcal{X}_{\overline{F}}$ is a family of stacky curves on $\mathcal{X}_{\overline{F}}$ which dominates $\mathcal{X}_{\overline{F}}$. One defines an intersection number between a family of movable curves and $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, as the generic value of the intersection of the members of the family with $(L,x)$ [DY24, Definition 8.7]. Motivated by the fact that the effective cone is the dual of the cone of movable curves [BDPP13], one defines $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathcal{X})$ to be the subset of $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$ which intersect non-negatively every movable family of stacky curves on $\mathcal{X}_{\overline{F}}$. If a height is defined by $(L,x)\in\operatorname{NS}_{\operatorname{orb}}(\mathcal{X})$, one defines
@@
Finally, we need one more definition before stating the main conjecture: given a stack $\mathcal{Y}$ and a ring $R$, denote by $\mathcal{Y}\langle R\rangle$ the set of $R$-isomorphism classes of $R$-points of $\mathcal{Y}$.
@@ 1.2. Results
The main theorem of this paper proves the Conjecture 1.1 on several modular curves. For each $N\in\mathbb{Z}_{>0}$, let $\mathscr{X}_{0}(N)$ be the compactified Deligne–Rapoport modular stack of elliptic curves endowed with a cyclic rational $N$-isogeny over $F$ (see Definition 3.2 for the precise definition). Consider
@@
Then the coarse moduli space $X_{0}(N)$ of $\mathscr{X}_{0}(N)$ is geometrically rationally connected if and only if it has genus 0, that is if and only if $N\in\mathfrak{N}_{0}$ (c.f. [Maz78]). So we assume that $N\in\mathfrak{N}_{0}$. The $F$-points of these stacks of bounded naive height have been counted by the papers [Bru92, Gra00, PPV20, PS21, Phi22, Mol23, MV23, BS24, AHPP25]. When $N=1$, the naive height on $\mathscr{X}_{0}(1)\langle F\rangle\cong\mathcal{P}(4,6)\langle F\rangle\to\mathbb{R}_{\geq 0},$ is the height function as in Definition 3.9. For example, when $F=\mathbb{Q}$ and the $\mathbb{Q}$-elliptic curve given by $Y^{2}=X^{3}+eX+f,$ where $e,f\in\mathbb{Z}$ and for all $p$ prime, $p^{4}\nmid e$ or $p^{6}\nmid f$, the naive height is given by $\max(|e^{3}|,|f^{2}|)$. (Note that some sources use a different normalization, e. g. in [AHPP25], it is given by $\max(4|e^{3}|,27|f^{2}|),$ but it has no effect on $a$- and $b$-invariants.) The naive height on $\mathscr{X}_{0}(N)\langle F\rangle$ refers to the composite function $H:\mathscr{X}_{0}(N)\langle F\rangle\to\mathscr{X}_{0}(1)\langle F\rangle\to\mathbb{R}_{>0}$, where $\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ forgets the level structure. By Lemma 3.4, one has that $\mathscr{X}_{0}(N)\langle F\rangle$ is Zariski-dense in $\mathscr{X}_{0}(N)$.
@@
Let $N\in\mathfrak{N}_{0}$ as above and let $M_{\operatorname{naive}}\in\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ be the element corresponding to the naive height $H$. When $F=\mathbb{Q}$, Conjecture 1.1 holds for $\mathscr{X}_{0}(N)$ with the respect to the height $H$.
@@
For every $N$ and $M_{\operatorname{naive}}$ from Theorem 1.2, one has the following values of $a$- and $b$-invariants:
@@
To prove Theorem 1.3, we first give a concrete description of $\mathscr{X}_{0}(N)$ as a square root stack over a stacky curve. Recall that given a line bundle $L$ on a stack $\mathcal{X}$, the $n$-th root stack of $L$, denoted by $\sqrt[n]{L/\mathcal{X}}$, is a stack defined in [Cad07], endowed with a morphism to $\mathcal{X}$ such that the pullback of $L$ is $n$-th power in the Picard group of $\sqrt[n]{L/\mathcal{X}}$.
The square root stacks are $\mu_{2}$-gerbes: they are the trivial $\mu_{2}$-gerbes if and only if $L$ is a square in the Picard group $\operatorname{Pic}(\mathcal{X})$. The square root stacks are classified by $\operatorname{Pic}(X)/2\operatorname{Pic}(X)$. In [AOV08], one describes a notion of rigidification, which essentially kills a subgroup scheme of generic stabilizer group. Hence if $\mathcal{X}$ is a stacky curve and $L$ line bundle, the generic stabilizer is $\mu_{2}$ and the rigidification of $\sqrt[2]{L/\mathcal{X}}$ is just $\mathcal{X}$. We denote the rigidification of $\mathcal{Y}$ by $\mathcal{Y}^{\operatorname{rig}}$.
Similarly, to notion of root stack of a line bundle, there is a notion of $(n_{1},\dots,n_{r})$-th root stack of an $r$-tuple of Cartier divisors $C_{1},\dots,C_{r}$ on an algebraic stack $\mathcal{X}$. When the dimension of $\mathcal{X}$ is equal to $1$ and $C_{1},\dotsc,C_{r}$ are mutually disjoint, then the root stack $\sqrt[n_{1},\dots,n_{r}]{C_{1},\dots,C_{r}/\mathcal{X}}\to\mathcal{X}$ can be thought as adding stackiness $\mu_{n_{i}}$ along $C_{i}$.
We have a canonical representable morphism $J_{N}:\mathscr{X}_{0}(N)\to\mathscr{X}_{0}(1)$ and an induced representable morphism $J_{N}^{\operatorname{rig}}:\mathscr{X}_{0}(N)^{\operatorname{rig}}\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. Let $\Lambda$ be the Hodge line bundle on $\mathscr{X}_{0}(1)$. There exists a unique line bundle $\Lambda^{\operatorname{rig}}$ on $\mathscr{X}_{0}(1)^{\operatorname{rig}}$ such that $\Lambda^{\otimes 2}$ is isomorphic to the pullback of $\Lambda^{\operatorname{rig}}$ under the rigidification morphism $\mathscr{X}_{0}(1)\to\mathscr{X}_{0}(1)^{\operatorname{rig}}$. We set $\Lambda_{N}:=J_{N}^{*}\Lambda$ and $\Lambda_{N}^{\operatorname{rig}}:=(J_{N}^{\operatorname{rig}})^{*}\Lambda^{\operatorname{rig}}$.
Let $K_{6}$ and $K_{4}$ be the divisors $X^{2}+3Y^{2}=0$ and $X^{2}+Y^{2}=0$, respectively, on $\mathbb{P}^{1}$. Let $E_{0},E_{1728}$ be the divisors corresponding to $X=0$ and $X=1728Y$, respectively.
@@ 1.4. Organization
In Section 2, we recall cyclotomic stacks, stacky Proj construction, root stacks, and sectors of DM stacks. Section 3 starts with the definition and geometric properties of modular curves $\mathscr{X}_{0}(N)$. After that, the naive height of $\mathscr{X}_{0}(N)$ is studied in Section 3.1, which concludes with the description of the element defining the naive height in $\operatorname{NS}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ (Corollary 3.13). In Section 3.2, we consider the orbifold canonical divisor $K_{\mathscr{X}_{0}(N),\operatorname{orb}}$ and the anticanonical naive height. In Section 3.3, we classify adequate cases and then prove Theorem 1.3 for these cases. Section 4.1 considers local computations needed to understand stacky curves over $\mathscr{X}_{0}(N)$, which is a preparation for the partial description of the orbifold pseudoeffective cone $\operatorname{\overline{Eff}}_{\operatorname{orb}}(\mathscr{X}_{0}(N))$ described in Section 4.2. In the remaining subsections of Section 4, Theorem 1.3 is proved for the non-adequate cases.
@@ 2.1. Cyclotomic stacks and the stacky Proj construction
Given a $\mathbb{Q}$-stack $\mathcal{X}$, [AHPP25, Lemma 2.7] explains necessary and sufficient conditions for $\mathcal{X}$ to be a stacky Proj, which involves definitions from [AH11, §2]; so let us summarize those definitions over $\operatorname{Spec}F$. A separated $F$-stack $\mathcal{X}$ of locally finitely presentation over $\operatorname{Spec}F$ is called a cyclotomic stack if for every geometric point $\overline{x}$ of $X$, the stabilizer group space $\operatorname{Aut}_{\mathcal{X}}(\overline{x})$ (or $\operatorname{Aut}(\overline{x})$ for short) is isomorphic to $(\mu_{n})_{\kappa(\overline{x})}$ for some $n\geq 1$. A line bundle $\mathcal{L}$ on a cyclotomic $F$-stack $\mathcal{X}$ is uniformizing if the associated classifying morphism $\mathcal{X}\to B\mathbb{G}_{m}$ is representable; equivalently, for every geometric point $\overline{x}$ of $\mathcal{X}$, $\operatorname{Aut}(\overline{x})$-action on $\mathcal{L}_{\overline{x}}$ is faithful (c.f. [AH11, Proposition 2.3.10]). Finally, a uniformizing line bundle $\mathcal{L}$ on a proper cyclotomic $F$-stack $\mathcal{X}$ is polarizing if there exists $m\geq 1$ such that $\mathcal{L}^{m}\cong c^{*}M$, where $c:\mathcal{X}\to X$ is the coarse moduli map and $M$ is an ample line bundle on $X$. A mild generalization of [AHPP25, Lemma 2.7] is as follows:
@@ 2.2. Root stacks
Given an effective Cartier divisor $\mathcal{D}$ on $\mathcal{X}$, pick a global section $s_{\mathcal{D}}$ of $\mathcal{O}_{\mathcal{X}}(\mathcal{D})$ that cuts out $\mathcal{D}$. Then the $n^{\rm th}$ root stack of $\mathcal{D}$ on $\mathcal{X}$ is defined as the fiber product $\sqrt[n]{\mathcal{D}/\mathcal{X}}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\mathcal{X}\times_{[\mathbb{A}^{1}/\mathbb{G}_{m}]}[\mathbb{A}^{1}/\mathbb{G}_{m}]$ of $\mathcal{X}\to[\mathbb{A}^{1}/\mathbb{G}_{m}]$ corresponding to $(\mathcal{L},s_{\mathcal{D}})$ and the $n^{\rm th}$-power morphism $\mathbb{A}^{1}/\mathbb{G}_{m}]\to[\mathbb{A}^{1}/\mathbb{G}_{m}]$; the isomorphism class of $\sqrt[n]{\mathcal{L}/\mathcal{X}}$ is independent of the choice of $s_{\mathcal{D}}$. Similarly, given a fixed choice $s_{\mathcal{D}}$, a $T$-object of $\sqrt[n]{\mathcal{D}/\mathcal{X}}$ is instead a quadruple $(x,M,u,\varphi)$ where $u\in H^{0}(M)$ and $\varphi(u^{n})=x^{*}s_{\mathcal{D}}$. Finally, when $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ are mutually disjoint effective Cartier divisors of $\mathcal{X}$, then $\sqrt[n_{1},\dotsc,n_{m}]{\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}/\mathcal{X}}$ is the iterative root stack $\sqrt[n_{1}]{\mathcal{D}_{1}/\sqrt[n_{2}]{\mathcal{D}_{2}/\dotsb/\sqrt[n_{m}]{\mathcal{D}_{m}/\mathcal{X}}}}$; note that the isomorphism class of this stack remains the same under permutation of $(n_{1},\mathcal{D}_{1}),\dotsc,(n_{m},\mathcal{D}_{m})$. Because our stacks of interests are DM stacks of dimension one, the mutually disjoint condition of divisors $\mathcal{D}_{1},\dotsc,\mathcal{D}_{m}$ is not restrictive.
@@
Observe that the rigidification morphism $\pi\colon\mathcal{X}\to\mathcal{X}^{\operatorname{rig}}$ is a $\mu_{n}$-gerbe by definition. Because $\pi^{*}\mathcal{L}^{\operatorname{rig}}\cong\mathcal{L}^{n}$, this isomorphism induces a morphism $h\colon\mathcal{X}\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}$ of $\mu_{n}$-gerbes over $\mathcal{X}^{\operatorname{rig}}$. Then for every smooth $F$-covering $U\to\mathcal{X}^{\operatorname{rig}}$ by an $F$-scheme $U$, the pullback $h_{U}\colon\mathcal{X}\times_{\mathcal{X}^{\operatorname{rig}}}U\to\sqrt[n]{\mathcal{L}^{\operatorname{rig}}/\mathcal{X}^{\operatorname{rig}}}\times_{\mathcal{X}^{\operatorname{rig}}}U$ is an isomorphism by [Ols16, Lemma 12.2.4]. Therefore, $h$ is also an isomorphism by descent. ∎